OpenStudy (lacedupcountry):

I have a math/Science question, i just need someone to check my answers! i will give away medals, i just feel like im not doing it right. I'll post the question below!

OpenStudy (lacedupcountry):

Here is my table

OpenStudy (lacedupcountry):

Data Analysis: Calculate the percentage of the cell cycle spent in each stage. Number of cells in given stage ÷ total number of cells counted × 100 = % of the cell cycle spent in this stage Here are my answers: Part 1: Interphase: 34 x 100 /49= 69.38% Prophase: 8*100/13=61.53% Metaphase: 3*100/4= 75% Anaphase: 2*100/3=66.67% Telophase: 2*100/4=50% Cytokinesis: 1*100/3=33.33% Part 2: Interphase: 49*100/34=144.11% Prophase: 13*100/8=162.5% Metaphase: 4*100/3=133.33% Anaphase: 3*100/2=150% Telophase: 4*100/2=200% Cytokinesis:3*100/1=300%

OpenStudy (anonymous):

do you need to round?

OpenStudy (lacedupcountry):

It doesn't say to in the directions

OpenStudy (anonymous):

oh ok, just to be sure i would round to hundredth like you did... part 1: interphase - 69.39 prophase - 61.54 part 2: interphase - 144.12

OpenStudy (anonymous):

https://www.youtube.com/watch?v=gCYcHz2k5x0

OpenStudy (anonymous):

the rest are correct

OpenStudy (anonymous):

in example, 56.179 <- the 9 is higher than 5, therefore you round 7 to 8. if the 9 were a 5,6,7, or 8, it would also make 7 an 8. If the 9 would have been less than 5, then the 7 would stay.

OpenStudy (anonymous):

got it?

OpenStudy (lacedupcountry):

Yes thanks!

OpenStudy (anonymous):

welcome